OpenStudy (anonymous):

Can someone help me change –5x + 2y = 12 into slope-intercept form.

OpenStudy (anonymous):

solve the equation for y

OpenStudy (anonymous):

so 2y = 12 + 5x and y = 6 + (5/2)x which is y = (5/2)x + 6

OpenStudy (anonymous):

im not really good at solving equations, honestly im not

OpenStudy (anonymous):

slope intercept = y=mx+b So you solve so it looks like that.